OpenStudy (lanhikari22):

integral(ln x dx) from [0,1]. Normally. you'd get negative infinity for that, right? ( Because 0 - infinity is - infinity ), or |-inf| = inf, since some say that area should be absolute. Anyway, is there any way to play with the form as to get -1 instead?

OpenStudy (anonymous):

Sorry, I misread it the first time. The value of that integral is -1, what's your question?

OpenStudy (anonymous):

\[ \int_0^1 \ln(x) dx = \left[ x\ln(x)-x\right]^1_0 = -1 - 0 = -1 \]

OpenStudy (anonymous):

\[\int\limits \ln (x)*1 dx=\ln x \int\limits 1 dx-\int\limits \frac{ d }{ dx }\left( \ln x \right)\int\limits 1dx=(\ln x)x-\int\limits \frac{ 1 }{ x }*x dx\] \[=x \ln x-x\]

OpenStudy (lanhikari22):

I guess my question is why 0*(ln 0) - 0 = 0 doesn't ln 0 approach negative infinity? 0 * - inf = what? Isn't that undetermined and dependent on which is faster anyhow? Or does it really just simply reduce to zero?

jimthompson5910 (jim_thompson5910):

@LanHikari22 the issue occurs when you reach the point `0*infinity` that's an indeterminate form (one of many) https://en.wikipedia.org/wiki/Indeterminate_form

OpenStudy (anonymous):

Ah, yeah. It turns out that \(x\) approaches zero "faster" than \(\ln(x)\) approaches negative infinity, and the product goes to zero. This can be demonstrated in a number of ways, for example via L'Hospital's Rule.

jimthompson5910 (jim_thompson5910):

you'll have to rewrite \[\Large x\ln(x)\] into \[\Large \frac{\ln(x)}{1/x}\] then use L'Hospital's rule

OpenStudy (lanhikari22):

That just takes you into an infinite deriving loop though due to it all being indefinite

OpenStudy (anonymous):

No it doesn't. Applying L'Hospital's rule gives \[\frac{1/x}{-1/x^2} =-x \] which approaches zero as \(x\rightarrow 0\).

OpenStudy (lanhikari22):

Oh crap. For whatever reason, I never thought of cancelling, I just took myself into an infinite loop. This all makes sense now, thanks!